OpenStudy (anonymous):

If cos2 x + 2cosx - 2 = 0, then cosx = _____.

OpenStudy (solomonzelman):

cos²x+2cosx-2=0 , like this ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh my bad

OpenStudy (anonymous):

in this case its still a quadratic though.

OpenStudy (solomonzelman):

Okay, so \(\normalsize\color{blue}{ \cos^2 x + 2\cos x - 2 = 0}\) :) For better vision let cos(x)=a, you come out with, \(\normalsize\color{blue}{ a^2 + 2a - 2 = 0}\) lets complete the square. \(\normalsize\color{blue}{ a^2 + 2a - 2 = 0}\) \(\normalsize\color{blue}{ a^2 + 2a +1 = 3}\) \(\normalsize\color{blue}{ (a+1)^2= 3}\) \(\normalsize\color{blue}{ a= -1±\sqrt{3}}\)

OpenStudy (anonymous):

oh!!! thank you!!

OpenStudy (solomonzelman):

yes phys, it is just a quadratic, technically, besides that for 'a' we can have 1 or 2 solutions, and for cosine function there are infinite number of solutions in many different intervals.

OpenStudy (solomonzelman):

You welcome, @garcia22and

OpenStudy (anonymous):

could you help me with another problem? please

OpenStudy (solomonzelman):

yes, perhaps.

OpenStudy (anonymous):

In triangle ABC, a = 12, angleB = 25°, and angleC = 45°. Find b.

OpenStudy (anonymous):

sin(110)/12 = sin(25)/b

OpenStudy (anonymous):

I mean A) 5.4 B) 6.6 C) 7.2 D) 22

OpenStudy (solomonzelman):

No phys, angle A is not given yet, although, \(\normalsize\color{blue}{ ∠~A~~=~~180-(~~~∠~B + ∠~C~~)}\) .

OpenStudy (solomonzelman):

You can't use the sine law yet, but after you find angle a, please go for it.

OpenStudy (anonymous):

Sin(A)/a = Sin(B)/b = Sin(C)/c

OpenStudy (anonymous):

so you solve for angle A. 180-25-45 = ?

OpenStudy (anonymous):

\[\frac{ \sin110 }{ 12 }= \frac{ sinB }{ b }\]

OpenStudy (anonymous):

right and we have Sin B

OpenStudy (solomonzelman):

yes, we have the angle B, it is 25

OpenStudy (anonymous):

when I get here what do I do next\[bsin110=5.071\]

OpenStudy (anonymous):

to make it simpler you could just say 12sin(25)/sin(110) = B

OpenStudy (anonymous):

I divide sin110?

OpenStudy (anonymous):

think about it algebraicly

OpenStudy (anonymous):

we want to isolate B

OpenStudy (anonymous):

ok I got it, thanks!

OpenStudy (anonymous):

Answer is 5.4

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

cool!! another problem: Which of the following functions has a period of 4π ? A) y = sinx B) y = sinx C) y = sin2x D) y = sin4x I thought it was B

OpenStudy (anonymous):

so for sin functions the period = 2pi/B

OpenStudy (anonymous):

uh for answers a and b are the same?

OpenStudy (anonymous):

oh sorry A is y=sin1/4x and B is sin1/2x

OpenStudy (anonymous):

2pi/B = 4pi (period) solve for B get 1/2, so the answer is sin1/2x

OpenStudy (anonymous):

ah okay so I was right, cool!

OpenStudy (anonymous):

The vertical component of a force of 100 pounds that makes an angle of 30° with the horizontal is 50 pounds. Is this true?

OpenStudy (anonymous):